# Self-Tuning: Instructing LLMs to Effectively Acquire New Knowledge through Self-Teaching

### v3

@@ 3 Wiki-Newpages-2023-QA: Datasets for Studying LLM Knowledge Acquisition
3.1 Document Collection
Given the well-structured nature of Wikipedia articles, which encompass extensive factual information and cover a wide range of topics across various domains, we gather documents from Wikipedia NewPages222https://en.wikipedia.org/wiki/Special:NewPages. This collection includes new articles from diverse domains published after the pre-training cut-off time of the LLMs being evaluated, allowing us to largely ensure that the models have not been exposed to these facts. To construct the document corpus, we specifically gather articles from September to October 2023, resulting in a total of 4,257 articles.333The pre-training cut-off for the Llama2 family models used in this study is 2022. Following Jiang et al. (2024b), we only utilize the first paragraph of each article, which provides a comprehensive summary and contains a wealth of factual information.
3.2 QA Pair Generation
To gather QA pairs, we utilize GPT-4 OpenAI (2023) along with our manually curated prompts to generate a variety of questions and their corresponding answers, aiming to cover all factual information within the given document. Specifically, we construct QA datasets for the open-ended generation and NLI tasks by employing the prompts shown in Table 18 and Table 19, respectively. A simplified example document with associated QA pairs is provided in Table 2. More detailed examples can be found in Appendix F.
@@ 3.2 QA Pair Generation
3.3 Splitting
To enable a comprehensive analysis in single-domain, multi-domain, and cross-domain situations, we develop three datasets and divide them into training and testing subsets.
@@ 3.3 Splitting
We partition the Wiki-Bio and Wiki-Multi datasets, comprising the document corpus and the derived QA datasets, into training and testing subsets for conducting evaluations in single-domain and multi-domain contexts. We directly utilize the Wiki-Film dataset as the test set for the cross-domain scenario. It is crucial to note that the training QA datasets only contain the QA pairs from open-ended generation tasks, ensuring a fair comparison with existing knowledge injection approaches. We provide extensive statistical information for the three datasets in Table 1 and a thorough analysis of the QA types in Appendix H.
4 Self-Tuning
4.1 Overview
Our objective is to equip an LLM , parameterized by , with the ability to learn how to derive knowledge from raw documents. This is achieved by training the model using a combination of training document dataset and associated training QA dataset , as depicted in the upper left part of Figure 2. To enhance effective knowledge absorption, we present along with a series of knowledge-intensive tasks (a.k.a. self-teaching tasks) that are related to their content for Self-Teaching (in the lower part of Figure 2). These tasks are generated in a self-supervised manner based on the contents of , using the proposed Self-Teaching learning approach (Section 4.2). The multi-task training objective is presented as follows:
Stage 2: Learn New Knowledge while Reviewing QA Skills.
Our aim is to train the model to apply the learned strategy for spontaneously extracting new knowledge from unseen documents (i.e., the test document dataset ). In addition to training on , we include , allowing the model to review and refine its question-answering ability. This approach enhances knowledge extraction from . The training objective is as follows:
Stage 3: Continually Learn.
Our goal is to ensure that the model thoroughly absorbs the new knowledge from unseen documents by conducting follow-up training on the dataset. The objective for this stage is as follows:
4.2 Self-Teaching Learning Strategy
Memorization.
Comprehension.
Gist identification improves the model’s ability to pinpoint the key elements (i.e., entities) within the atomic facts. Specifically, we prompt the model with Highlight the key information within the article:, and use the entities within the document as gold answers, identified using Spacy444https://spacy.io/usage.
Natural language inference provides the model with the capability to determine whether a statement can be inferred from specific document contents (i.e., “Yes,” “No,” or “It’s impossible to say”), thus avoiding misconceptions that may arise during knowledge acquisition. Specifically, we use a randomly sampled sentence (identified using NLTK555A natural language toolkit. https://www.nltk.org/) within the document content as the true statement, and a corrupted version where one entity is replaced by an irrelevant entity from another sentence as the false statement. Then, we prompt the model with Based on the article above can we conclude that and the sampled sentence (either initial or corrupted), with the three relations as options and corresponding answers.
Self-Reflection.
5 Experiments
In this section, we assess the efficacy of Self-Tuning in enhancing an LLM’s ability to acquire new knowledge on the curated Wiki-Newpages-2023-QA datasets in three distinct scenarios: single-domain, multi-domain, and cross-domain.
5.1 Setup
@@
| Standard Instruction-Tuning | train doc&test doc | train QA | |
@@
| Self-Tuning via Read. | train QA&train doc (reading-comprehension format) | test doc | |
@@
Datasets and Evaluation Metrics. We evaluate the effectiveness of Self-Tuning in both knowledge acquisition and retention for a well-rounded analysis, considering that catastrophic forgetting Wang et al. (2023) frequently occurs during knowledge acquisition.
(1) Evaluation on Knowledge Acquisition We carry out assessments on three knowledge acquisition tasks, namely memorization, extraction, and reasoning. For memorization, we utilize test document datasets and report perplexity (PPL) Jelinek et al. (1977). For extraction, we employ test QA datasets for open-ended generation tasks and evaluate factual accuracy using exact match (EM), Recall, F1 Kwiatkowski et al. (2019), and Rouge-L Lin (2004); Jiang et al. (2024b). We also assess accuracy using the bidirectional entailment approach with the Deberta-Large-MNLI model He et al. (2021). For reasoning, we use test QA datasets for NLI tasks and report accuracy.
(2) Evaluation on Knowledge Retention We carry out assessments on two aspects of knowledge retention, specifically extraction and reasoning. For extraction, we evaluate the model’s performance in retaining factual knowledge using Natural Questions (NQ) Kwiatkowski et al. (2019) (NQ-open Min et al. (2021)) and report EM and F1 scores. For reasoning, we assess the capability in retaining commonsense knowledge using CommonsenseQA (CSQA) Talmor et al. (2019) and report accuracy.
All evaluations are conducted in a closed-book setting. Details can be found in Appendix J.
Compared Methods. We compare our method with the following representative approaches, as presented in the upper part of Table 7 and report the mean results of three different runs.
Continued Pre-training trains the model on the dataset.
Standard Instruction-tuning first trains on both and datasets, then fine-tunes on dataset.
PIT Jiang et al. (2024b) first trains on and datasets to equip the model with the ability to absorb knowledge from raw documents, with the QA pairs placed right before the corresponding document texts, then trains on the data.
Implementation Details.
We utilize Llama2-7B and Llama2-13B for our investigation and provide an analysis on Llama2-7B-chat in Appendix C for a comprehensive understanding. We use the following training objectives: for training on document data , we compute the standard next-token prediction loss by averaging over all tokens in the document (Equation 4); for training on QA data , we compute the average negative log-likelihood loss only on tokens in the answer given the question (Equation 5), where and refer to the length of the tokenized document sequence and answer sequence, respectively. Full details can be found in the Appendix K.
5.2 Main Results
| Memorization | Extraction | Reason. | Extraction | Reasoning | ||||||
| Llama2-7B | ||||||||||
| w/o Knowledge Injection | ||||||||||
| w/ Knowledge Injection | ||||||||||
| Con. Pre-training | 7.28 | 6.33 | 3.62 | 15.96 | 18.72 | 16.11 | 15.09 | 16.00 | 24.11 | 53.40 |
| Llama2-13B | ||||||||||
| w/o Knowledge Injection | ||||||||||
| w/ Knowledge Injection | ||||||||||
Table 4 shows the evaluation results on Llama2-7B and Llama2-13B w.r.t., knowledge acquisition and retention in the single-domain scenario using the Wiki-Bio dataset. We make the following observations:
The curated dataset has minimal overlap with the LLMs’ pre-training data. The extremely poor performance in the closed-book setting (e.g., with EM around 2% for knowledge extraction) suggests that the dataset shares little similarity with the pre-training data, thereby making the evaluation results reliable. The non-zero EM values can be attributed to a small number of the gathered Wikipedia articles that were initially published but underwent revisions after the cut-off time.
Continued pre-training for knowledge acquisition proves challenging across all three dimensions. We find that continuing pre-training on new documents may result in a decline in knowledge extraction performance (on Llama2-13B), compared to the baseline performance. This could be due to the fact that merely continuing pre-training might adversely affect its question-answering capability, even when equipped with new knowledge, as demonstrated by the lowered PPL. This observation is consistent with the findings in Cheng et al. (2024). Moreover, the marginal improvements in memorization (reducing PPL by 1-2%) and reasoning (increasing accuracy by 2-7%) suggest that such a naive approach fails to help the model memorize and capitalize on new knowledge. This highlights the importance of evaluating the model’s knowledge acquisition ability comprehensively across multiple dimensions.
Self-Tuning substantially improves the LLM’s knowledge acquisition ability. Self-Tuning greatly enhances the performance of Llama2-7B and Llama2-13B across three dimensions: reducing PPL to nearly 1, signifying effective memorization of the new documents; increasing EM by roughly 20% on the knowledge extraction task, attaining performance comparable to the open-book setting; achieving high accuracy among the compared methods for the reasoning task, demonstrating excellent understanding of the newly acquired knowledge. These results underscore the value of comprehension and self-reflection, beyond simply memorizing document contents. This confirms the effectiveness of the Self-Teaching learning approach.
Self-Tuning excels in knowledge retention. Unlike other methods that display fluctuating performance in knowledge retention, Self-Tuning demonstrates a strong ability to maintain previously acquired knowledge in terms of both knowledge extraction and reasoning. The slight improvements in evaluation metrics, such as F1 (roughly 1-3% on extracting learned world knowledge) and accuracy (approximately 5-13% on commonsense reasoning), compared to the closed-book performance without knowledge injection, suggest that systematically learning new knowledge doesn’t necessarily lead to catastrophic forgetting. Instead, it enhances the elicitation and understanding of previously learned knowledge.
5.3 Training Dynamics
The remarkable performance of Self-Tuning in enhancing knowledge acquisition does not stem from early-fitting. In Figure 3, we observe that Self-Tuning not only memorizes new knowledge more rapidly than the compared methods, lowering PPL to almost 1 within 3 epochs, but also consistently achieves the best performance during long-term training. Remarkably, Self-Tuning begins to outperform the open-book performance from the 5th epoch and reaches its peak at the 25th epoch with a 5% higher EM score on the knowledge extraction task. This observation highlights the importance and potential of incorporating knowledge into the parameters of LLMs.
Self-Tuning performs well in preserving previously acquired knowledge, with only a small decline in EM of approximately 2-3% over the course of 50 training epochs. This suggests that Self-Tuning has great potential for real-world applications.
5.4 Results in the Multi-Domain and Cross-Domain Scenarios
| Method | Acquisition | NQ (Reten.) | CSQA (Reten.) | |||||||
| Memorization | Extraction | Reason. | Extraction | Reasoning | ||||||
| w/o Knowledge Injection | ||||||||||
| w/ Knowledge Injection | ||||||||||
| w/o Knowledge Injection | ||||||||||
| w/ Knowledge Injection | ||||||||||
| PIT | 1.52 | 6.39 | 2.67 | 16.97 | 18.92 | 17.10 | 3.03 | 13.06 | 23.42 | 54.38 |
To explore the potential of Self-Tuning for enhancing LLM’s knowledge acquisition and retention in real-world scenarios, we evaluate its performance in two challenging situations (Table 5): the multi-domain scenario (in the upper part), where both the training documents and test documents come from various complex domains; the cross-domain scenario (in the lower part), where training data and test documents belong to entirely different domains, i.e., the training data is from Wiki-Bio, while the test data is from Wiki-Film.
Self-Tuning consistently shows strong potential in enhancing knowledge acquisition and retention in both settings. In Table 5, Self-Tuning consistently achieves the best performance in both settings, suggesting the potential to expand this method to a wider range of documents containing diverse new knowledge.
The capacity to systematically absorb knowledge improves generalization ability. The substantial improvements over all compared methods in the cross-domain setting, e.g., exceeding EM by 13% on the knowledge extraction task, highlight the importance of equipping the model with the ability to systematically absorb knowledge from raw documents using the Self-Teaching strategy, rather than solely teaching it how to answer questions.
5.5 Variants of Self-Tuning
| Mem. | Extraction | Reason. | Extraction | Reasoning | ||||||
| w/ Knowledge Injection | ||||||||||
[…]
Setup. To further investigate the effectiveness of Self-Tuning, we present three variations, as depicted in Table 8: (1) Self-Tuning w/o Review, where we continue training on test documents without the reviewing capability; (2) Self-Tuning via Read., which displays the training documents in a reading-comprehension format Cheng et al. (2024) (an example is shown in Table 22); (3) Self-Tuning w/ Pre-Review, which trains on a combination of training documents and training QA in the second stage, before training on test documents.
Results. As depicted in Table 5.5, despite having lower performance than Self-Tuning, all variations significantly enhance the model’s ability for knowledge acquisition compared to simply continued pre-training, which further validates the effectiveness of Self-Tuning in improving knowledge acquisition.
Reviewing the QA ability aids in knowledge acquisition and retention. Compared to Self-Tuning, Self-Tuning w/o Review also displays inferior performance on the knowledge retention task. Moreover, we suspect that the slightly lower performance of Self-Tuning w/ Pre-Review is due to reviewing the QA ability during the continuous learning of new knowledge helps in reducing the distribution shift, thereby stabilizing the training process. These findings underscore the importance of reviewing the QA ability during the continuous knowledge acquisition.
5.6 Ablation Study
5.7 Fine-grained Comparison
@@
Results. As shown in Table 7, we observe that Self-Tuning consistently outperforms PIT in the overall evaluation and the fine-grained evaluations related to different QA types. These findings underscore the importance of equipping the model with the ability to systematically acquire new knowledge. Furthermore, we present a qualitative comparison between the answers generated by PIT and Self-Tuning in Appendix D. To gain insights into potential enhancements for Self-Tuning, we also conduct a detailed error analysis on the types of factual errors that remain challenging after implementing Self-Tuning in Appendix E.
6 Conclusion
In this study, we introduce Self-Tuning, a framework aimed at enhancing an LLM’s ability to effectively learn from raw documents through self-teaching. Specifically, we create Self-Teaching, a self-supervised learning strategy that presents documents as plain texts along with various knowledge-intensive tasks derived directly from the documents, without the need for any mining patterns. Furthermore, we present three Wikipedia-Newpages-2023-QA datasets to enable a comprehensive evaluation of an LLM’s knowledge acquisition capabilities, w.r.t., memorization, extraction, and reasoning across three distinct scenarios. Our findings show that Self-Tuning consistently yields superior performance on the knowledge acquisition tasks while demonstrating impressive knowledge retention performance. These results suggest the potential for broader applications of the proposed Self-Tuning, such as acquiring domain-specific knowledge.
Limitations
Conducting Experiments on Various LLMs.
Due to constraints in time and computational resources, our extensive experiments are conducted on Llama2 family models. For future research, we plan to explore the effectiveness of our approach in enhancing the knowledge acquisition capabilities of other models such as Mistral-7B Jiang et al. (2023), Orca2-7B Mitra et al. (2023), Llama3 models MetaAI (2024), and larger-scale models like Llama2-70B.
Applying to Broader Scenarios.
Our study primarily centers on infusing new factual knowledge into the parameters of LLMs to keep them up-to-date. However, our proposed Self-Teaching strategy doesn’t necessitate any mining patterns to build the knowledge-intensive tasks based on the new corpus. As a result, we foresee that our Self-Tuning framework can be utilized in various areas, such as arming the LLMs with domain-specific knowledge Cheng et al. (2024); Que et al. (2024), and mathematical principles Xu et al. (2024).
Appendix A Knowledge Editing for Knowledge Injection
Knowledge editing Mitchell et al. (2022); Zheng et al. (2023); Yao et al. (2023); Jiang et al. (2024a); Zhang et al. (2024a) is also recognized as one related research direction in the field of knowledge injection. It concentrates on rectifying outdated or inaccurate factual knowledge stored in the model, without affecting other facts. In contrast, our focus lies in enabling LLMs to efficiently acquire knowledge from raw corpora.
| Method | Training Data in Each Stage | ||
| Stage 1 | Stage 2 | Stage 3 | |
| Continued Pre-training | test doc }x5 | ||
| Standard Ins.-tuning | train doc& test doc }x3 | train QA }x1 | |
| PIT | train QAtrain doc }x3 | test doc }x3 | |
| Self-Tuning | train QA&train doc w/ self-teaching tasks }x2 | train QA&test doc }x1 | test doc }x2 |
| Variants of Self-Tuning | |||
| Self-Tuning w/o Review | train QA&train doc w/ self-teaching tasks }x2 | test doc }x3 | |
| Self-Tuning via Read. | train QA&train doc (reading-comp. format) }x2 | test doc }x3 | |
| Self-Tuning w/ Pre-Review | train QA&train doc w/ self-teaching tasks }x2 | train QA&train doc }x1 | test doc }x3 |
| Additional Compared Methods | |||
| Standard Ins.-Tuning w/o Forget. | train doc& test doc }x3 | train QA & test doc }x1 | |
| PIT++ | train QA }x1 | train QAtrain doc }x3 | test doc }x3 |
| Mixed Training | train doc&train QA& test doc }x3 | ||
| Method | Wiki-Newpages-2023-10-Bio (Acquisition) | NQ (Reten.) | CSQA (Reten.) | ||||||||
| Mem. | Extraction | Reason. | Extraction | Reasoning | |||||||
| Llama2-7B | |||||||||||
| w/ Knowledge Injection | |||||||||||
| Continued Pre-training | 7.28 | 4.68 | 2.87 | 14.63 | 16.98 | 15.07 | 7.96 | 16.05 | 24.67 | 53.40 | |
[…]
| Mixed Training | 1.42 | 24.13 | 20.67 | 38.82 | 39.95 | 38.66 | 55.69 | 19.33 | 28.40 | 58.97 | - |
Appendix B Evaluation Results of Additional Compared Methods
For a thorough assessment, we also examine the efficiency of our suggested Self-Tuning by contrasting it with three other notable methods, namely standard instruction-tuning without forgetting, PIT++, and mixed training, as displayed in Table 8. We present the evaluation results in Table A. Our Self-Tuning consistently shows superior performance, for example, it increases the EM by 11% on the knowledge extraction task.
Appendix C Evaluation Results on Llama2-7B-chat in the Single-domain Scenario
| Memorization | Extraction | Reason. | Extraction | Reasoning | ||||||
| Llama2-7B-chat | ||||||||||
| w/o Knowledge Injection | ||||||||||
| w/ Knowledge Injection | ||||||||||
Appendix D Qualitative Analysis

### v4

@@ Appendix B Wiki-Newpages-2023-QA: Datasets for Studying LLM Knowledge Acquisition
B.1 Document Collection
Given the well-structured nature of Wikipedia articles, which encompass extensive factual information and cover a wide range of topics across various domains, we gather documents from Wikipedia NewPages666https://en.wikipedia.org/wiki/Special:NewPages. This collection includes new articles from diverse domains published after the pre-training cut-off time of the LLMs being evaluated, allowing us to largely ensure that the models have not been exposed to these facts. To construct the document corpus, we specifically gather articles from September to October 2023, resulting in a total of 4,257 articles.777The pre-training cut-off for the Llama2 family models used in this study is 2022. Following Jiang et al. (2024c), we only utilize the first paragraph of each article, which provides a comprehensive summary and sufficient factual information.
B.2 QA Pair Generation
To gather QA pairs, we utilize GPT-4 OpenAI (2023) along with our manually curated prompts to generate a variety of questions and their corresponding answers, aiming to cover all factual information within the given document. Note that GPT-4 is solely used to convert document knowledge into QA pairs without introducing any additional information. It can be replaced by any other model.
Specifically, we construct QA datasets for the open-ended generation and NLI tasks by employing the prompts shown in Table 23 and Table 24, respectively. A simplified example document with associated QA pairs is provided in Table 6. More detailed examples can be found in Appendix O.
@@ B.2 QA Pair Generation
B.3 Splitting
To enable a comprehensive analysis in single-domain, multi-domain, and cross-domain situations, we develop three datasets and divide them into training and testing subsets, ensuring zero knowledge overlap.
@@ B.3 Splitting
We partition the Wiki-Bio and Wiki-Multi datasets, comprising the document corpus and the derived QA datasets, into training and testing subsets for conducting evaluations in single-domain and multi-domain contexts. We directly utilize the Wiki-Film dataset as the test set for the cross-domain scenario. It is crucial to note that the training QA datasets only contain the QA pairs from open-ended generation tasks, ensuring a fair comparison with existing knowledge injection approaches. We provide extensive statistical information for the three datasets in Table 1 and a thorough analysis of the QA types in Appendix Q.
@@
| Standard Ins.-tuning | train doc&test doc | train QA | |
@@
| Self-Tuning via Read. | train QA&train doc (reading-comp. format) | test doc | |
@@
| Additional Compared Methods | |||
| Standard Ins.-Tuning w/o Forget. | train doc&test doc | train QA&test doc | |
| PIT++ | train QA | train QAtrain doc | test doc |
| Mixed Training | train doc&train QA&test doc | ||
| Method | Wiki-Newpages-2023-QA (Acquisition) | NQ (Reten.) | CSQA (Reten.) | ||||||||
| Mem. | Extraction | Reason. | Extraction | Reasoning | |||||||
| Knowledge Acquisition on Wiki-Newpages-2023-10-Bio (Single-Domain Scenario) | |||||||||||
| w/o Knowledge Injection | |||||||||||
| Open-book w/ test doc | 8.41 | 55.20 | 31.83 | 64.48 | 75.55 | 62.10 | 7.96 | - | - | - | |
| w/ Knowledge Injection | |||||||||||
| Cont. Pre-training | 7.28 | 6.33 | 3.62 | 15.96 | 18.72 | 16.11 | 15.09 | 16.00 | 24.11 | 53.40 | |
| Standard Ins.-tuning | 6.83 | 6.94 | 5.13 | 19.15 | 19.05 | 19.48 | 39.09 | 15.72 | 23.67 | 51.84 | |
| Standard Ins.-Tuning w/o Forget. | 2.82 | 9.35 | 7.09 | 21.25 | 21.72 | 21.51 | 36.08 | 16.05 | 24.88 | 54.30 | |
| PIT | 2.08 | 14.03 | 11.61 | 27.15 | 28.86 | 27.11 | 11.93 | 15.72 | 26.31 | 57.58 | |
| Mixed Training | 1.42 | 24.13 | 20.67 | 38.82 | 39.95 | 38.66 | 55.69 | 19.33 | 28.40 | 58.97 | |
| Knowledge Acquisition on Wiki-Newpages-2023-10-Multi (Multi-Domain Scenario) | |||||||||||
| w/o Knowledge Injection | |||||||||||
| Open-book w/ test doc | 7.84 | 48.93 | 26.63 | 60.37 | 71.71 | 58.54 | 6.33 | - | - | - | |
| Closed-book | 7.84 | 4.53 | 2.73 | 16.19 | 18.63 | 16.38 | 6.33 | 16.05 | 24.67 | 53.40 | |
| w/ Knowledge Injection | |||||||||||
| Cont. Pre-training | 3.32 | 5.86 | 3.40 | 18.04 | 20.59 | 18.42 | 14.51 | 17.02 | 25.05 | 53.56 | |
| Standard Ins.-tuning | 2.73 | 8.66 | 5.73 | 24.94 | 25.64 | 25.31 | 34.91 | 15.60 | 26.26 | 52.74 | |
| PIT | 1.96 | 14.31 | 8.72 | 30.26 | 33.97 | 30.22 | 10.69 | 15.55 | 27.02 | 55.12 | |
| Self-Tuning | 1.13 | 22.30 | 16.51 | 39.94 | 41.02 | 39.89 | 50.65 | 16.34 | 25.85 | 69.29 | |
| Knowledge Acquisition on Wiki-Newpages-2023-(9)10-Film (Cross-Domain Scenario) | |||||||||||
| w/o Knowledge Injection | |||||||||||
| Open-book w/ film doc | 8.30 | 57.38 | 34.45 | 68.64 | 78.92 | 66.31 | 7.35 | - | - | - | |
| Closed-book | 8.30 | 3.35 | 1.88 | 11.27 | 12.97 | 11.49 | 7.35 | 16.05 | 24.67 | 53.40 | |
| w/ Knowledge Injection | |||||||||||
| Cont. Pre-training | 5.52 | 3.46 | 2.30 | 11.83 | 14.30 | 11.98 | 12.04 | 16.79 | 25.35 | 56.02 | |
| Standard Ins.-tuning | 2.83 | 5.23 | 3.77 | 16.15 | 17.45 | 16.45 | 51.69 | 14.41 | 25.54 | 49.80 | |
| PIT | 1.52 | 6.39 | 4.50 | 16.97 | 18.92 | 17.10 | 3.03 | 13.06 | 23.42 | 54.38 | |
| Self-Tuning | 1.10 | 22.51 | 16.44 | 35.58 | 36.60 | 35.43 | 44.92 | 16.77 | 26.44 | 66.34 | |
Appendix C Evaluation Results on Llama2-7B
To thoroughly assess our proposed Self-Tuning method, we compare its efficiency against three other notable approaches: standard instruction-tuning without forgetting, PIT++, and mixed training, as shown in Table 9. The evaluation results, presented in Table 8, demonstrate that Self-Tuning consistently outperforms the alternatives. For instance, it improves EM by 11% in the knowledge extraction task.
Combining the results from Table 8 with the training strategies in Table 9, we emphasize the importance of first training the model to develop the ability to absorb new knowledge before training on test documents. This finding aligns with the conclusions of Jiang et al. (2024c).
Notably, Self-Tuning enables the model to absorb new knowledge from incoming test documents more efficiently. Unlike mixed training—which requires retraining on both training documents, training QA, and test documents—Self-Tuning leverages the capabilities acquired in the initial training stage to directly learn from test documents, needing only a review of QA ability. This makes it significantly more training-efficient over time.
Appendix D Fine-grained Comparison
@@
Results. As shown in Table 9, we observe that Self-Tuning consistently outperforms PIT in the overall evaluation and the fine-grained evaluations related to different QA types. These findings underscore the importance of equipping the model with the ability to systematically acquire new knowledge. Furthermore, we present a qualitative comparison between the answers generated by PIT and Self-Tuning in Appendix E. To gain insights into potential enhancements for Self-Tuning, we also conduct a detailed error analysis on the types of factual errors that remain challenging after implementing Self-Tuning in Appendix G.
Appendix E Qualitative Analysis